# Towards Effective Prompt Stealing Attack against Text-to-Image Diffusion Models

### v1

@@ A-A VLMasExpert
A-B User Study
We consider several important factors for the user study. First, we conduct our investigation in as vague a way as possible to alleviate potential safety threats. Specifically, to avoid publicity of stealing technology, we don’t inform the volunteers what our task is. In terms of the detailed question design, we ask about the similarity between the two images (i.e., the showcase and recovered image), specifying no identity of what each image represents. Second, we make much effort to eliminate bias. On the one hand, for each question, we disrupt the method order randomly. This can avoid habitual choice and ensure the independence of each question. On the other hand, we survey a wide range of people with various backgrounds, including gender, age, and education. Among all the respondents, 57.28% are men while 42.72% are women. As for age, 41.75% of the surveyed people were between 18-25, and 45.63% of those were between 26-30. Including respondents aged between 31-40 (total proportion of 10.68%), these three parts account for 98.06% of the total. Of the remaining 1.94%, the people aged under 18 account for 0.97% (one sample), and those between 51-60 account for 0.97% (one sample). We also realize the effect of education on respondents’ preferences. We keep all the questions to 8th grade or lower to ensure they are fully understood by the respondents. Among all the surveyed people, almost all of them obtained a college degree or above, and only 0.97% of them (one sample) own a high school degree. Specifically, 3.88% of the total respondents have a college degree, and a proportion of 30.1% have a bachelor’s degree. The left part, with a proportion of 65.05%, occupies the vast majority, and owns a postgraduate degree. We refer to PromptStealer [43] for option design.
@@ A-B User Study
A-C Defense
@@
PromptShield. PromptShield proposes to defend prompt-stealing attacks by turning the showcase into adversarial examples. They set the ground-truth prompt without artist modifiers as the target adversarial prompt. Then the defender uses the gradient of the modifier predictor to optimize the showcase. There are three defects in this defense. First, PromptShield assumes white-box access to the attacker’s modifier predictor, which may not be practical. Second, they assume that all the prompts sold on the prompt market have artist modifiers so that they can construct the target adversarial prompts. However, this is not practical for all prompts. For example, the prompt shown in Figure 1, which is sold on PromptBase, has no artist modifier. Third, if there is an artist’s style in the prompt, removing this element from the showcase may damage its attractiveness.
@@
Detail of defenses. We set the hyperparameter of conventional defenses following the principle that the perturbation should not affect human perception of the subjects and styles. Specifically, for random noise, we generate Gaussian noise with a mean of 0 and a standard deviation of 25, which is almost perceptible to humans. For the puzzle effect, we add a 4×4 grid to the showcase, with variability set to 3. We generate a text watermark “@watermark” as the intellectual property claim. The text font size is set to 20, and the gaps in the rows and columns are set to 20 and 30, respectively. We evaluate the defense performance on the RealPrompt dataset to simulate the preventive effect in the real world.
@@
Adaptive attack. The proposed defense, named PromptGuard, employs the gradient of a unified loss function to generate adversarial perturbations. The loss curve is shown in Figure 16, where the unified loss ultimately converges around 0.45, demonstrating the defense’s effectiveness. However, our observations reveal that the primary decline in the loss comes from BLIP, while CLIP contributes minimally. This can be attributed to CLIP’s pre-training on LAION [40], a dataset containing billions of examples, which makes it inherently robust to adversarial perturbations. In contrast, BLIP is more fragile due to its significantly smaller training dataset. Based on these findings, when designing adaptive attacks, we focus on modifying the BLIP structure, as adversarial examples exhibit poor transferability across different BLIP models.

### v2

@@ B-A VLMasExpert
B-B User Study
We specify the user study setting as follows:
@@ B-B User Study
Study Detail.
The recruited volunteers were asked to complete a questionnaire on perceptual image similarity. Each participant viewed 12 randomly selected groups of images (a showcase and its recovered versions); each group contained 6 image pairs (5 baselines and our method), yielding 72 questions in total. All images were non-explicit, stylistic paintings within the range of everyday visual content (the examples can be seen in Figure 6). For each question, volunteers rated the perceptual similarity between the showcase and the recovered image using the same scoring criteria as PromptStealer [68]. To avoid publicizing the prompt-stealing setting, we did not disclose the precise research goal and only told participants they were comparing the similarity between two images. To reduce potential bias, we anonymized the methods for every question and randomized the method order within each group to prevent habitual choices and ensure independence across questions.
Volunteers.
We totally recruited 103 volunteers and obtained 103 valid questionnaires. To eliminate bias, the volunteers are recruited from various backgrounds, including gender, age, and education. Specifically, among all the respondents, 57.28% are men while 42.72% are women. As for age, 41.75% of the surveyed people were between 18-25, and 45.63% of those were between 26-30. Including respondents aged between 31-40 (total proportion of 10.68%), these three parts account for 98.06% of the total. Of the remaining 1.94%, both the people aged under 18 and those between 51-60 account for 0.97%. We also realize the effect of education on respondents’ preferences. We keep all the questions to 8th grade or lower to ensure they are fully understood by the respondents. Among all the surveyed people, most of them obtained a college degree or above, and only 0.97% of them have a high school degree. Specifically, 3.88% of the total respondents have a college degree, and a proportion of 30.1% have a bachelor’s degree. The left part, with a proportion of 65.05%, occupies the vast majority, and owns postgraduate degree.
@@
B-C Defense
PromptShield.
PromptShield seeks to defend against prompt-stealing by turning showcases into adversarial examples using gradients from the attacker’s modifier-predictor. It sets the ground-truth prompt without artist modifiers as the adversarial target and then optimizes the showcase with the predictor’s gradients. This approach has four shortcomings. First, it assumes the attacker trains a modifier classifier, which our attack does not, making PromptShield inapplicable and limiting its real-world utility. Second, it assumes white-box access to the attacker’s predictor, which is often unrealistic. Third, it presumes all marketplace prompts include artist modifiers so a target adversarial prompt can be constructed; however, many do not. For example, the prompt in Figure 1 from PromptBase lacks any artist modifier. Finally, even when a prompt contains an artist modifier, removing that element safeguards only the artist modifier while leaving other modifiers unprotected.
Detail of Defenses.
@@ B-C Defense
We set the hyperparameter of conventional defenses following the principle that the perturbation should not affect human perception of the subjects and styles. Specifically, for random noise, we generate Gaussian noise with a mean of 0 and a standard deviation of 25, which is almost perceptible to humans. For the puzzle effect, we add a 4×4 grid to the showcase, with variability set to 3. We generate a text watermark “@watermark” as the intellectual property claim. The text font size is set to 20, and the gaps in the rows and columns are set to 20 and 30, respectively. We evaluate the defense performance on the RealPrompt dataset to simulate the preventive effect in the real world.
Analysis of Adaptive Attack.
@@
The proposed defense, named PromptGuard, employs the gradient of a unified loss function to generate adversarial perturbations. The loss curve is shown in Figure 16, where the unified loss ultimately converges around 0.45, demonstrating the defense’s effectiveness. However, our observations reveal that the primary decline in the loss comes from BLIP, while CLIP contributes minimally. This can be attributed to CLIP’s pre-training on LAION [64], a dataset containing billions of examples, which makes it inherently robust to adversarial perturbations. In contrast, BLIP is more fragile due to its significantly smaller training dataset. Based on these findings, when designing adaptive attacks, we focus on modifying the BLIP structure, as adversarial examples exhibit poor transferability across different BLIP models. The core of defending against prompt-stealing attacks is to break the connection between the showcase and its caption, independent of the specific captioning model. Therefore, future defenses should prioritize mitigation strategies that remain robust across different attack model architectures.
B-D More Analysis
Main results.
[…]
Appendix C More Related Works
C-A Related Works on Large Language Models
There have been several prompt-stealing attacks against large language models. Yang et al. [81] propose a prompt attention algorithm to infer the detailed intent of prompts, which realizes accurate prompt stealing even from a single input-output pair. Hui et al. [19] propose to use a gradient-based method for optimizing the adversarial prompt to steal the system prompt. Sha et al. [66] introduce a parameter extractor and a prompt reconstructor to extract the prompt, and Tan et al. [75] propose a text gradient-based method to steal prompts in the wild iteratively. Nguyen et al. [42] propose to adaptively stack ciphers to jailbreak large reasoning models. There are also efforts to protect the privacy of LLMs. For example, Hao et al. [12] initiate the study of private inference on Transformer-based models, and propose the first systematic ZK proof framework for non-linear mathematical functions using the perspective of table lookup [11].
C-B Related Works on T2I Diffusion Models
Attacks
. Given the popularity and great commercial value of T2I models and prompts, lots of work has been dedicated to studying their vulnerability. Wu et al. [79] use a prompt, captioned with a query image, to conduct a membership inference attack on black-box T2I models. Following this, several works also study the risk of membership leakage [8, 16]. Carlini et al. [5] propose a data extraction attack against T2I models via querying the training prompt. Lukas et al. [74] inject backdoors into the text encoder for predefined image generation. Huang et al. [18] utilize personalization strategies (a trigger in the prompt) as a shortcut to realize a few-shot backdoor, and Liang et al. [32] propose a dual-embedding guided backdoor attack against multimodal contrastive learning. Then Liang et al. [30] propose a multimodal attribution backdoor attack to improve the attack generalizability. Recent research focuses on the jailbreak vulnerability [40], where attackers intentionally modify the prompt, inducing the model to generate offensive content even when the safety filter is applied [54, 57, 82]. For instance, Natalie et al. [38] propose the concept of adversarial prompts and design a black-box framework for attacking image generation models. Li et al. [25] propose a scalable red teaming method via distribution modeling for text-to-image generative systems. Zhao et al. [85] explore the vulnerability of the memory mechanism and introduce a multiturn jailbreak attack against the text-to-image generation system. Liu et al. [35] first propose to jailbreak text-to-video models with discrete prompt optimization.
Defenses
. As for defenses, Wu et al. [80] propose to regulate personalization models in embedding, and Li et al. [28] propose a comprehensive framework to protect data in the generative AI era. Guo et al. [10] propose a copyright defense method that integrates poisoned sample detection and adaptive optimization training. Jing et al. [21] propose a text-driven soft prompt-guided defense framework to tackle unsafe generation. Liang et al. [31] introduce a model-agnostic jailbreak defense for text-to-video models. Liang et al. [29] first propose a formal and effective autonomous defense system for mobile agents. Zhao et al. [86] propose using benign triggers to counter data-poisoning backdoor attacks in LLMs.